OpenStudy (anonymous):

How many quarts of pure antifreeze must be added to 6 quarts of a 20% antifreeze solution to obtain a 60% antifreeze solution?

OpenStudy (amistre64):

amount @ 1 = 6(.2-.6)/(1-.6) maybe ...

OpenStudy (amistre64):

nope; gotta redo that one :)

OpenStudy (anonymous):

60% of 6 + x is x + 20% of 6 as an equation this is \[.6(x+6)=x+.2\times 6\]

OpenStudy (anonymous):

right hand side of the equation says you have x amount of pure antifreeze plus the 20% of the 6 you already have. left hand side says you want the resulting amount to be 60% antifreeze. now you have to solve for x

OpenStudy (anonymous):

i would multiply by 10 first but i am lazy.

OpenStudy (dumbcow):

lol work it out, x= 6 avg of 100 and 20 is 60 so it makes sense the num of quarts is equal

OpenStudy (anonymous):

\[6(x+6)=10x+2\times 6\] \[6x+36=10x+12\] \[36-12=10x-6x\] \[24=4x\] \[6=x\]

OpenStudy (anonymous):

lot of "6s" in this problem

OpenStudy (anonymous):

i hate math

OpenStudy (amistre64):

a+b = t ax+by = tz ax = (a+b)z - by ax = az +bz -by a(x-z) = b(z-y) a = b(z-y)/(x-z) where b = known amount, and y = its %, z = final%; and x = unknown amounts % a = 6(.6-.2)/(1-.6)

OpenStudy (dumbcow):

haha...

OpenStudy (anonymous):

i like dumbcow's answer though very snappy!

OpenStudy (anonymous):

so put 6?

OpenStudy (amistre64):

6 seems to be the consensus 6 + 6(.2) = 12(.6) 6 + 1.2 = 7.2 7.2 = 7.2 its good

OpenStudy (amistre64):

my first bout gave me -6 ... i wonder if that a side effect tho ...

OpenStudy (anonymous):

it was 6.. thanks guys! this is awesome.